# ME-IGM: Individual-Global-Max in Maximum Entropy Multi-Agent Reinforcement Learning

### v3

@@ Abstract
1 Introduction
@@
Figure 1: The figure illustrates the improvement of our approach compared to existing maximum entropy MARL methods. The left figure shows a straightforward approach to applying maximum entropy MARL in the CTDE context, where blue texts represent the desired objectives, and black texts indicate corresponding constraints. It reveals that existing maximum entropy MARL methods under the CTDE framework implicitly constrain the global Q-value to be the sum of local Q-values (as in VDN), significantly limiting the expressiveness of the critic network. The right figure depicts the improvements in ME-IGM. ME-IGM first applies any credit assignment mechanism that satisfies the IGM condition, such as QMIX and QPLEX, to obtain $Q_{i}$ for each agent $i$. Given the meaningful order of local Q-values (for different actions), an order-preserving transformation $f_{i}$ converts $Q_{i}$ to $f_{i}(Q_{i})$ as the policy logits. This transformation is optimized using a loss function that minimizes the expected difference between $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$, which guarantees monotonic policy improvement in maximum entropy MARL.
@@
Collaborative multi-agent tasks, where a team of agents works together to complete a task and receives joint rewards, are inherently challenging for reinforcement learning (RL) agents. This difficulty arises due to the absence of individual reward signals for each agent and the significantly higher exploration requirements compared to single-agent scenarios. The challenge is further exacerbated by the exponential growth of the joint action space as the number of agents increases. In this work, we propose leveraging maximum entropy RL to promote exploration and adopting the individual-global-max (IGM) credit assignment mechanism (Rashid et al. 2020) to effectively distribute joint rewards among the agents.
@@
Maximum entropy RL (Levine 2018), recognized for promoting exploration (Haarnoja et al. 2018), smoothing the objective landscape (Ahmed et al. 2019), and improving robustness (Eysenbach and Levine 2021), has been extensively applied in single-agent RL. However, as illustrated in Figure 1, naively extending maximum entropy RL to multi-agent settings under the centralized training and decentralized execution (CTDE) paradigm leads to uniform credit assignment, which limits the ability of multi-agent reinforcement learning (MARL) agents to learn from joint rewards.
@@
On the other hand, numerous studies have explored the development of effective credit assignment mechanisms. Among them, the IGM condition has demonstrated strong empirical performance by aligning the maximum local Q-values with the maximum global Q-value. Previous works have primarily applied the IGM condition in deterministic policy settings, where policies are derived using the argmax of local Q-values. However, in maximum entropy MARL methods, stochastic local policies are learned, and misalignment can occur between local policies and the maximum global Q-values. That is, selecting joint actions based on the highest logits of local policies does not necessarily lead to maximizing the global Q-value. This limitation highlights a critical challenge in achieving effective credit assignment within the maximum entropy MARL framework. This challenge, as we will demonstrate, is not unique to IGM-based methods but also represents a broader obstacle for actor-critic-based maximum entropy MARL algorithms, such as FOP (Zhang et al. 2021), hindering their ability to achieve optimal coordination.
@@
  • We confirm the effectiveness of our algorithm through extensive evaluation on matrix games, SMAC-v2 (Ellis et al. 2022), and Overcooked (Carroll et al. 2019), where our method attains state-of-the-art results.
2 Background
2.1 Multi-Agent Reinforcement Learning
Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek and Amato 2016). Formally, a Dec-POMDP is represented as a tuple $(\mathcal{A},S,U,T,r,O,G,\gamma)$, where $\mathcal{A}\equiv\{1,...,n\}$ is the set of $n$ agents. $S$, $U$, $O$, and $\gamma$ are state space, action space, observation space, and discount factor, respectively.
@@ 2.1 Multi-Agent Reinforcement Learning
2.2 Individual-Global-Max Condition
Individual-Global-Max (IGM) (Rashid et al. 2020) is a commonly used credit assignment method in value-decomposition-based cooperative MARL and is defined as follows:
Definition 2.1.
@@ 2.2 Individual-Global-Max Condition
Then, we say that $[Q_{i}]_{i=1}^{n}$ satisfy IGM for $Q_{tot}$ under $s_{t}$.
2.3 Related Works
@@
Multi-Agent Credit Assignment: CTDE MARL primarily consists of two algorithmic branches: actor-critic methods (Lowe et al. 2017; Yu et al. 2022; Lin et al. 2023; Foerster et al. 2018; Kuba et al. 2021; Zhong et al. 2023) and value decomposition methods (Sunehag et al. 2017; Rashid et al. 2020; Wang et al. 2020). Actor-critic methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence between local policies and the softmax of the global Q-value. On the other hand, value decomposition methods address the credit assignment problem by decomposing the global value function into multiple local value functions while satisfying the IGM condition. In particular, VDN (Sunehag et al. 2017) assumes that the global Q-function is the sum of local Q-functions. QMIX (Rashid et al. 2020) allows the mixing function, which aggregates local Q-values into the global Q-value, to be any function with non-negative partial derivatives, i.e., $\frac{\partial Q_{tot}}{\partial Q_{i}}\geq 0,\forall i\in\mathcal{A}.$ QTRAN (Son et al. 2019) introduces an additional set of inequality constraints to construct its loss function. QPLEX (Wang et al. 2020) uses the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize. Our method, ME-IGM, enhances the exploration capability of these MARL algorithms by leveraging the maximum entropy RL framework and is compatible with all the aforementioned value decomposition methods.
Maximum Entropy MARL: Maximum entropy RL is demonstrated to encourage exploration in single-agent RL (Haarnoja et al. 2018). Most maximum entropy MARL works utilize an actor-critic architecture, minimizing the KL divergence between the local actors and the softmax of the centralized critic (He et al. 2022; Guo and Wu 2022; Zhang et al. 2021). However, as demonstrated later, these methods cannot guarantee that the joint action (at each state) determined by local policies maximizes the global Q-value, thus violating the IGM condition. In this case, we propose a novel order preserving transformation to ensure that the IGM condition is met while employing maximum entropy MARL.
3 Misalignment Between Local Policies and the Maximum Global Q-Value
@@
The Misalignment Problem of Maximum Entropy MARL: While the IGM condition defines the relationship between local Q-values and the global Q-value, what is often more important in practice is the alignment between local policies and the global Q-value. In other words, joint actions that maximize the global Q-value can be obtained by querying local policies. Previous work in maximum entropy MARL (Zhang et al. 2021; Guo and Wu 2022) often decompose the global Q-value into individual local Q-values to satisfy the IGM condition. These local Q-values are then used to guide the training of local stochastic policies. Compared with value decomposition methods such as QMIX, they train stochastic policy networks in addition to the Q-functions, which can potentially improve multi-agent exploration in complex environments. However, while the value decomposition ensures alignment between global and local Q-values via the IGM condition, there is typically no explicit constraints imposed between the local policies and the local Q-values during the policy training phase. This lack of coupling often leads to misalignment during testing, where the local policies fail to consistently select joint actions that maximize the global Q-value, defeating the purpose of enforcing the IGM condition. Consequently, the learned policies may struggle to achieve optimal global coordination during execution.
@@
Figure 2: Illustration of misalignment between local policies and the maximum global Q-value. When naively combining the IGM condition with maximum entropy MARL, local policies often select suboptimal joint actions, leading to a lower Q-value, depicted by the blue curve. In contrast, the optimal joint action should achieve the global-max Q-value, shown as the orange curve.
@@
Empirical Demonstration: To analyze the performance gap caused by the misalignment between local policies and the maximum global Q-value, we evaluate the difference between the maximum global Q-value and the global Q-value selected by local policies in the SMAC-v2 zerg 5vs5 scenario, using a state-of-the-art Maximum Entropy MARL algorithm (Zhang et al. 2021). Figure 2 shows the results, averaged over five random seeds with standard deviations included. We can observe that there is a significant Q-value gap, emphasizing the potential for improvement by developing algorithms that enable the selection of joint actions with the highest global Q-value through local stochastic policies.
4 Order-Preserving Transformation
@@ 4.1 Formulation of OPT
Here, $W^{1}_{{}_{(d_{1}\times 1)}}$, $W_{{}_{(1\times d_{1})}}^{2}$, $b^{1}_{{}_{(d_{1}\times 1)}}$, and $b^{2}_{{}_{(1\times 1)}}$ are generated by a hyper-network [^3] For details of the hyper-network, please refer to the Appendix. which takes the global state $s$ as input. Consider $x$ as a local Q-value of dimension $(1\times d_{2})$, then $W^{1}x+b^{1}$ is a set of hidden variables of size $(d_{1}\times d_{2})$. Notably, when all entries in $W^{1}$ are non-negative, the relative order of elements in $x$ is preserved in each of the $d_{1}$ hidden variables. Next, a nonlinear activation function $\sigma$, such as ELU (Clevert 2015), is applied, followed by another order-preserving linear transformation using $W^{2}$ and $b^{2}$. Consequently, the final output, $\text{OPT}(x,s)$, maintains both the dimension and the relative order of the entries in the input $x$.
@@
4.2 Training Scheme of OPT
Figure 3: The overall pipeline of ME-IGM.
@@ 4.2 Training Scheme of OPT
There is a gap between $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})$ and $Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)$. Also, we find that the policy improvement using Equation (4) no longer guarantees a monotonic increase in the global Q-value, unlike the approach based on Equation (3). Specifically, we have the following theorem:
Theorem 4.1.
Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$. Updating the policy with Equation (3) ensures monotonic improvement in the global Q-value. That is, $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}.$
@@
11:   calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq. 12
@@
where $\gamma$ is the discount factor, $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function and $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})\triangleq|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}),s_{t+1})|$.
Proof is provided in the Appendix. This theorem implies that we must enforce $\epsilon\leq 0$ to guarantee monotonic policy improvement when using Equation (4) as the objective. This motivates us to minimize the factor $C$ and adopt the following objective to train the OPTs.
@@
$\pi_{jt}^{new}=(\pi_{1}^{new},\cdots,\pi_{n}^{new})$ and $\pi_{i}^{new}$ is defined in Eq. (5).

### v4

@@ Abstract.
Key words and phrases:
Multi Agent Reinforcement Learning; Maximum Entropy Reinforcement Learning; Centralized Training with Decentralized Execution
doi:
GYYC3346
1. Introduction
@@
Figure 1. The figure illustrates the improvement of our approach compared to existing maximum entropy MARL methods. The left figure shows a straightforward approach to applying maximum entropy MARL in the CTDE context, where blue texts represent the desired objectives, and black texts indicate corresponding constraints. It reveals that existing maximum entropy MARL methods under the CTDE framework implicitly constrain the global Q-value to be the sum of local Q-values (as in VDN), significantly limiting the expressiveness of the critic network. The right figure depicts the improvements in ME-IGM. ME-IGM first applies any credit assignment mechanism that satisfies the IGM condition, such as QMIX and QPLEX, to obtain $Q_{i}$ for each agent $i$. Given the meaningful order of local Q-values (for different actions), an order-preserving transformation $f_{i}$ converts $Q_{i}$ to $f_{i}(Q_{i})$ as the policy logits. This transformation is optimized using a loss function that minimizes the expected difference between $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$, which guarantees monotonic policy improvement in maximum entropy MARL.
[…]
This image demonstrates improvement of our approach compared to existing maximum entropy MARL methods
@@
Collaborative multi-agent tasks, where a team of agents works together to complete a task and receives joint rewards, are inherently challenging for reinforcement learning (RL) agents. This difficulty arises due to the absence of individual reward signals for each agent and the significantly higher exploration requirements compared to single-agent scenarios. The challenge is further exacerbated by the exponential growth of the joint action space as the number of agents increases. In this work, we propose leveraging maximum entropy RL to promote exploration and adopting the individual-global-max (IGM) credit assignment mechanism (rashid2020monotonic) to effectively distribute joint rewards among the agents.
@@
Maximum entropy RL levine2018reinforcement, recognized for promoting exploration haarnoja2018soft, smoothing the objective landscape ahmed2019understanding, and improving robustness eysenbach2021maximum, has been extensively applied in single-agent RL. However, as illustrated in Figure 1, naively extending maximum entropy RL to multi-agent settings under the centralized training and decentralized execution (CTDE) paradigm leads to uniform credit assignment, which limits the ability of multi-agent reinforcement learning (MARL) agents to learn from joint rewards.
@@
On the other hand, numerous studies have explored the development of effective credit assignment mechanisms. Among them, the IGM condition has demonstrated strong empirical performance by aligning the maximum local Q-values with the maximum global Q-value. Previous works have primarily applied the IGM condition in deterministic policy settings, where policies are derived using the argmax of local Q-values. However, in maximum entropy MARL methods, stochastic local policies are learned, and misalignment can occur between local policies and the maximum global Q-values. That is, selecting joint actions based on the highest logits of local policies does not necessarily lead to maximizing the global Q-value. This limitation highlights a critical challenge in achieving effective credit assignment within the maximum entropy MARL framework. This challenge, as we will demonstrate, is not unique to IGM-based methods but also represents a broader obstacle for actor-critic-based maximum entropy MARL algorithms, such as FOP (zhang2021fop), hindering their ability to achieve optimal coordination.
We propose ME-IGM, the first CTDE maximum entropy MARL method that is compatible with the IGM condition. In particular, to address the misalignment between stochastic local policies and the maximum global Q-values, we introduce an order-preserving transformation (OPT) that maps local Q-values to policy logits while preserving their relative order. This ensures that selecting actions with the highest logits collectively leads to a joint action that maximizes the global Q-value. Additionally, the OPT is trained with a theoretically grounded objective, guaranteeing monotonic policy improvement in CTDE MARL.
@@
  • We confirm the effectiveness of our algorithm through extensive evaluation on matrix games, SMAC-v2 (ellis2022smacv2), and Overcooked carroll2019utility, where our method attains state-of-the-art results.  [^1] We have published our code at https://github.com/WentseChen/Soft-QMIX
2. Background
2.1. Multi-Agent Reinforcement Learning
Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) oliehoek2016concise. Formally, a Dec-POMDP is represented as a tuple $(\mathcal{A},S,U,T,r,O,G,\gamma)$, where $\mathcal{A}\equiv\{1,...,n\}$ is the set of $n$ agents. $S$, $U$, $O$, and $\gamma$ are state space, action space, observation space, and discount factor, respectively.
@@ 2.1. Multi-Agent Reinforcement Learning
2.2. Individual-Global-Max Condition
Individual-Global-Max (IGM) rashid2020monotonic is a commonly used credit assignment method in value-decomposition-based cooperative MARL and is defined as follows:
[…]
@@ 2.2. Individual-Global-Max Condition
Then, we say that $Q_{i}]_{i=1}^{n}$ satisfy IGM for $Q_{tot}$ under $s_{t}$.
2.3. Related Works
@@
Multi-Agent RL. In the CTDE framework, MARL primarily involves two common algorithmic approaches: policy gradient methods lowe2017multi; yu2022surprising; lin2023tizero; foerster2018counterfactual; kuba2021trust; zhong2023heterogeneousagent and value function decomposition methods sunehag2017value; rashid2020monotonic; wang2020qplex. Policy gradient methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence. On the other hand, Value function decomposition addresses the credit assignment problem by decomposing the global value function into multiple local value functions. VDN sunehag2017value assumes that the global Q-function is the sum of local Q-functions. QMIX rashid2020monotonic; peng2021facmac permits the mixer function to be any function with non-negative weights. QTRAN son2019qtran optimizes an additional pair of inequality constraints to construct a loss function. QPLEX wang2020qplex use the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize. The expressive power of the critic networks in these algorithms increases sequentially, but experimental results show that QMIX has better performance hu2021rethinking. Therefore, in this work, QMIX is used as the mechanism for credit assignment.
Maximum Entropy MARL. Maximum entropy RL levine2018reinforcement is demonstrated to have advantages such as encouraging exploration haarnoja2018soft and increasing robustness eysenbach2021maximum in single-agent RL scenarios. In maximum entropy MARL, most works utilize an actor-critic (AC) architecture and typically aim to minimize the KL divergence between the local actor and the centralized critic he2022multiagent; guo2022learning; pu2021decomposed; zhang2021fop.
Comparison with Actor-Critic Approaches. Most existing maximum entropy MARL methods (e.g., mSAC pu2021decomposed, FOP zhang2021fop, and HASAC liu2023maximum) rely on the Actor-Critic framework. Our value-based ME-IGM differs from them in three aspects. First, Order Preservation: AC-based methods may suffer from a misalignment where the order of actions preferred by the local actor does not match the local Q-function due to approximation errors. In contrast, our method uses Order-Preserving Transformations (OPT) to strictly enforce this alignment, ensuring the IGM condition is met. Second, Objective Function: AC methods typically minimize KL divergence to distill policies. Our method minimizes the MSE loss between logits, which has been shown to be superior in distillation contexts kim2021comparing. Third, Specific Baselines: FOP zhang2021fop employs a credit assignment mechanism similar to QPLEX but relies on the Individual-Global-Optimal (IGO) assumption. While FOP argues that locally optimizing policies reduces the joint policy gap, it does not explicitly guarantee alignment. HASAC liu2023maximum avoids the IGO assumption through sequential optimization but ignores the misalignment problem between stochastic policies and Q-values. Our work identifies this misalignment as a critical bottleneck and proposes OPT to resolve it, thereby achieving state-of-the-art performance.
3. Misalignment Between Local Policies and the Maximum Global Q-Value
@@
The Misalignment Problem of Maximum Entropy MARL: While the IGM condition defines the relationship between local Q-values and the global Q-value, what is often more important in practice is the alignment between local policies and the global Q-value. In other words, joint actions that maximize the global Q-value can be obtained by querying local policies. Previous work in maximum entropy MARL zhang2021fop; guo2022learning often decompose the global Q-value into individual local Q-values to satisfy the IGM condition. These local Q-values are then used to guide the training of local stochastic policies. Compared with value decomposition methods such as QMIX, they train stochastic policy networks in addition to the Q-functions, which can potentially improve multi-agent exploration in complex environments. However, while the value decomposition ensures alignment between global and local Q-values via the IGM condition, there is typically no explicit constraints imposed between the local policies and the local Q-values during the policy training phase. This lack of coupling often leads to misalignment during testing, where the local policies fail to consistently select joint actions that maximize the global Q-value, defeating the purpose of enforcing the IGM condition. Consequently, the learned policies may struggle to achieve optimal global coordination during execution.
@@
Figure 2. Illustration of misalignment between local policies and the maximum global Q-value. When naively combining the IGM condition with maximum entropy MARL, local policies often select suboptimal joint actions, leading to a lower Q-value, depicted by the blue curve. In contrast, the optimal joint action should achieve the global-max Q-value, shown as the orange curve.
[…]
Illustration of misalignment
@@
Empirical Demonstration: To analyze the performance gap caused by the misalignment between local policies and the maximum global Q-value, we evaluate the difference between the maximum global Q-value and the global Q-value selected by local policies in the SMAC-v2 zerg 5vs5 scenario, using a state-of-the-art Maximum Entropy MARL algorithm (zhang2021fop). Figure 2 shows the results, averaged over five random seeds with standard deviations included. We can observe that there is a significant Q-value gap, emphasizing the potential for improvement by developing algorithms that enable the selection of joint actions with the highest global Q-value through local stochastic policies.
4. Order-Preserving Transformation
@@ 4.1. Formulation of OPT
Here, $W^{1}_{{}_{(d_{1}\times 1)}}$, $W_{{}_{(1\times d_{1})}}^{2}$, $b^{1}_{{}_{(d_{1}\times 1)}}$, and $b^{2}_{{}_{(1\times 1)}}$ are generated by a hyper-network [^2] For details of the hyper-network, please refer to the Appendix. which takes the global state $s$ as input. Consider $x$ as a local Q-value of dimension $(1\times d_{2})$, then $W^{1}x+b^{1}$ is a set of hidden variables of size $(d_{1}\times d_{2})$. Notably, when all entries in $W^{1}$ are non-negative, the relative order of elements in $x$ is preserved in each of the $d_{1}$ hidden variables. Next, a nonlinear activation function $\sigma$, such as ELU clevert2015fast, is applied, followed by another order-preserving linear transformation using $W^{2}$ and $b^{2}$. Consequently, the final output, $\text{OPT}(x,s)$, maintains both the dimension and the relative order of the entries in the input $x$.
@@
4.2. Training Scheme of OPT
Figure 3. The overall pipeline of ME-IGM.
[…]
The overall pipeline of ME-IGM.
@@ 4.2. Training Scheme of OPT
There is a gap between $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})$ and $Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)$. Also, we find that the policy improvement using Equation (4) no longer guarantees a monotonic increase in the global Q-value, unlike the approach based on Equation (3). Specifically, we have the following theorem: {theorem} Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$. Updating the policy with Equation (3) ensures monotonic improvement in the global Q-value. That is, $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}.$
@@
11:   calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq.
@@
where $\gamma$ is the discount factor, $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function, $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})$, and $\Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})\triangleq|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})$- $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}),s_{t+1})|$.
The proof is provided in Appendix B, where we also analyze how this error affects the general convergence properties of policy improvement. This motivates us to minimize the factor $C$ and adopt the following objective to train the OPTs.
@@
where $\pi_{jt}^{new}=(\pi_{1}^{new},\cdots,\pi_{n}^{new})$ and $\pi_{i}^{new}$ is defined in Eq. (5).
Table 1. (a) presents the payoff matrix for a one-step matrix game involving two agents: the row player and the column player. Each agent has three possible actions: $\{\text{A, B, C}\}$. The rewards for each joint action are given in a $3\times 3$ matrix. This payoff structure is non-monotonic, as the optimal action for one player depends on the action chosen by its teammate. (b)–(d) illustrate the results of ME-QMIX, FOP, and QMIX, respectively. The first row and first column display the policies adopted by the row and column players after $10k$ training steps. The $3\times 3$ matrix in the bottom right corner represents the global Q-function $Q_{\text{tot}}$ learned by each algorithm. We can see that ME-QMIX is the only method that assigns the highest probability to selecting the optimal joint action, highlighting its effectiveness in overcoming the misalignment issue and achieving optimal coordination. $\epsilon$ in (d) denotes the exploration rate used in $\epsilon$-greedy exploration.